Circular No. 9041

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

NOVA SAGITTARII 2009 K. Kinugasa, Gunma Astronomical Observatory (GAO), writes that

K. Nishiyama (Kurume, Fukuoka, Japan) and F. Kabashima (Miyaki-cho, Saga, Japan) reported a brightening of a USNO-B1.0-catalogue star located at α = 17^{h}44^{m}08^{s}.478, δ = −26^{o}05′47′′.37 (equinox 2000.0) — from fainter than mag 13.5 on their unfiltered CCD patrol image (105-mm-f.l. camera lens) taken on Apr. 18.705 UT to mag 12.5 on Apr. 21.681 (with additional magnitudes as follows: Apr. 21.711, 11.7; 22.720, 11.7; 23.760, 12.3; 25.683, 12.2; 27.815, 12.4). Kinugasa, S. Honda, and O. Hashimoto took lowdispersion (range 400–800 nm, resolution ∼ 500) and high-dispersion (range 480–664 nm, resolution ∼ 40000) spectra of this object on Apr. 27.7 with the GAO 1.5-m telescope (+ GLOWS and GAOES, respectively), which show a broad (FWHM ∼ 2600 km/s) and double-peaked Hα emission, with several emission lines of Fe II and O I — indicating that this object is a classical nova in decline phase. The astrometry (using the USNO-A2.0 reference stars) with images also taken on Apr. 28.8 (when the nova was at V = 13.7, R = 12.9, I = 10.6) with the 1.5-m telescope (+ GLOWS), reveals position end figures 08^{s}.44, 48′′.7 — which is ∼ 1′′.4 from the USNO-B1.0 star, so the nova may be unrelated to it. H. Yamaoka, Kyushu University, adds that the ASAS-3 system (Pojmanski 2002, Acta Astr. 52, 397) also detects the object at the following V magnitudes: Apr. 16.374, [14.0:; 19.304, 12.4; 22.332, 13.4; 25.285, 13.6; 30.293, [14.0:.

SATELLITE OF SATURN The IAU Working Group for Planetary System Nomenclature has ap-

proved a new designation and name of a satellite of Saturn (cf. IAUC 8873):

Saturn LIII Aegaeon = S/2008 S 1  IAUC 9023

COMET C/2002 S7 = C/2008 N4 = C/1996 X3 (SOHO) Further to IAUC 8985 and R. Kracht’s suggestion that C/2002 S7 =

C/2008 N4, Kracht has now confirmed the likely identification with C/1996 X3 (cf. IAUC 8734). Since the 1996 comet passed perihelion 0.12 day later than the gravitational 2002–2008 linkage indicates, he assumed the nongravitational parameters A_{1} = 0.0000, A_{2} = +0.0027. Further details are given on MPEC 2009-J14, which includes the following orbital elements: T = 2008 July 4.38 TT, q = 0.0482, e = 0.9851, ω = 52^{o}.39, Ω = 49^{o}.82, i = 13^{o}.47 (equinox 2000.0), P = 5.78 yr.

2009 May 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green